OpenStudy (anonymous):

There are 98 students in 8 homeroom classes. What is the ratio of students to classes?

OpenStudy (anonymous):

98/8

OpenStudy (anonymous):

Simplify the following: 98/8 Reduce 98/8 to lowest terms. The gcd of 98 and 8 is 2, so 98/8 = (2×49)/(2×4) = 2/2×49/4 = 49/4: Answer: | | 49/4

OpenStudy (anonymous):

Thank you! you rock

OpenStudy (anonymous):

I know @Brookebrookiee